# Image-based Multimodal Models as Intruders: Transferable Multimodal Attacks on Video-based MLLMs

### v2

@@ Abstract
Video-based multimodal large language models (V-MLLMs) have shown vulnerability to adversarial examples in video-text multimodal tasks. However, the transferability of adversarial videos to unseen models—a common and practical real-world scenario—remains unexplored. In this paper, we pioneer an investigation into the transferability of adversarial video samples across V-MLLMs. We find that existing adversarial attack methods face significant limitations when applied in black-box settings for V-MLLMs, which we attribute to the following shortcomings: (1) lacking generalization in perturbing video features, (2) focusing only on sparse key-frames, and (3) failing to integrate multimodal information. To address these limitations and deepen the understanding of V-MLLM vulnerabilities in black-box scenarios, we introduce the Image-to-Video MLLM (I2V-MLLM) attack. In I2V-MLLM, we utilize an image-based multimodal model (IMM) as a surrogate model to craft adversarial video samples. Multimodal interactions and temporal information are integrated to disrupt video representations within the latent space, improving adversarial transferability. In addition, a perturbation propagation technique is introduced to handle different unknown frame sampling strategies. Experimental results demonstrate that our method can generate adversarial examples that exhibit strong transferability across different V-MLLMs on multiple video-text multimodal tasks. Compared to white-box attacks on these models, our black-box attacks (using BLIP-2 as surrogate model) achieve competitive performance, with average attack success rates of 55.48% on MSVD-QA and 58.26% on MSRVTT-QA for VideoQA tasks, respectively. Our code will be released upon acceptance.
@@ 1 Introduction
Figure 1: An example of transferable adversarial attack on different target V-MLLMs for VideoQA task.
@@
Recent work has shown that video-based multimodal large language models (V-MLLMs) are vulnerable to adversarial video samples [21], even though they have achieved remarkable performance on a wide range of video-text multimodal tasks [22, 18, 24, 29, 9, 42]. Existing work primarily focuses on white-box attacks, where information about the target model is accessible. However, the transferability of adversarial video samples across V-MLLMs remains unexplored, which is a more common and practical setting in real-world scenarios. It is still uncertain whether the adversarial videos generated on the source model can effectively attack other target models, posing significant security risks to the deployment of V-MLLMs in real-world applications.
@@
In this paper, we pioneer an investigation into the transferability of adversarial video samples across V-MLLMs. Through detailed analysis in Sec. 3.2, we think previous methods have these shortcomings: (1) lacking generalization in perturbing video features, (2) focusing only on sparse key-frames, and (3) failing to integrate multimodal information. FMM attack [21] is the first proposed white-box attack method targeting V-MLLMs. It utilizes flow-based temporal mask to select key-frames and applies perturbations to these frames. FMM attack performs well in the white-box setting but has limited transferability in the black-box setting. FMM attack heavily relies on the video features, which causes the generated perturbations to overfit to the video features extracted by the surrogate model, thereby reducing their generalizability. Additionally, since FMM attack applies perturbations only to key-frames, it cannot ensure that all frames sampled by the target model are perturbed. Taking low-level image features into account can help with improving transferability of adversarial samples. Previous image-to-video cross-modal attacks [39, 37, 19] demonstrate the possibility of using image models as surrogates to attack video models in the black-box setting. However, these traditional attack methods typically focus on the video classification tasks with vision-only models, failing to integrate multimodal information.
@@
To address these limitations, we propose a highly transferable attack method, named as Image To Video MLLM (I2V-MLLM) attack (see Fig. 2). In I2V-MLLM, we utilize an image-based multimodal model (IMM) as a surrogate model to craft adversarial video samples without accessing the internals of target V-MLLMs. Specifically, we extract key-frames from videos and send them into an IMM to obtain adversarial perturbations. Multimodal interactions and temporal information are integrated to disrupt video representations within the latent space, improving adversarial transferability. In addition, a perturbation propagation technique is introduced to handle different unknown frame sampling strategies used by V-MLLMs.
Figure 2: The overview of our proposed I2V-MLLM attack.
@@
We conduct various experiments on three well-established datasets, MSVD-QA [41], MSRVTT-QA [41], and ActivityNet-200 [15] to evaluate the performance of our proposed I2V-MLLM attack in multiple video-text multimodal tasks. The experimental results demonstrate that our method can generate adversarial videos with strong transferability across different V-MLLMs (Chat-Univi [18], LLava-Next-Video [44], VideoChat [22], Video-LLaMA [42]), and achieve competitive performance with white-box attacks against V-MLLMs. Our main contributions are summarized as follows:
@@
  • We explore the transferable adversarial attack on four different V-MLLMs and analyze the reasons for the low transferability when using existing methods to generate adversarial video samples (see Sec. 3.2). To the best of our knowledge, this is the first work to explore black-box attacks on V-MLLMs.
@@
  • We propose a highly transferable attack method, named I2V-MLLM, for V-MLLMs using IMMs to generate adversarial video samples (see Sec. 3.3). The adversarial videos generated by this method can effectively disrupt different V-MLLMs, significantly degrading their performance on multiple video-text multimodal tasks.
@@
  • We conduct extensive experiments on four different V-MLLMs using MSVD-QA, MSRVTT-QA, and ActivityNet-200 (see Sec. 4.2 and Sec. 4.3). The results demonstrate that our proposed attack method has strong transferability across V-MLLMs.
@@ 2.1 Multimodal large language models
MLLMs typically consist of a vision model, a pretrained LLM, and a projector that translates visual information into textual representations that the LLM can process. Currently, MLLMs can be categorized into image-based and video-based types. Image-based MLLMs [9, 26, 46, 1, 2, 16, 3] are designed to handle image-text inputs. They excel in tasks such as visual question answering, image captioning, and more. V-MLLMs extend the capabilities of image-based MLLMs by incorporating temporal modules that allow them to understand and process video input. This enables them to perform tasks like video question answering (VideoQA), spatio-temporal localization, and video captioning. For example, Chat-UniVi [18] extracts specific frames from videos and utilizes DPC-KNN [12] to group these frames into distinct events, Video-LLaMA [42] employs sequential encoding to capture temporal relationships among video frames, VideoChatGPT [29] applies temporal pooling to video features to extract temporal information. These methods equip the models with the capability to capture and interpret temporal dynamics, thus enabling a more comprehensive understanding of video content.
@@ 2.2 Adversarial attacks on MLLMs
Despite the impressive performance, MLLMs are highly susceptible to adversarial attacks [45, 28, 8, 36, 43, 4, 27]. For image-based MLLMs, several studies have assessed their vulnerabilities to adversarial attacks. Fu et al. [13] introduce Trojan-like images that force the target models to invoke malicious external tools or APIs specified by the attacker. Dong et al. [11] utilize open-source MLLMs to generate transferable adversarial examples capable of attacking closed-source commercial models like Bard [14], Bing Chat [31], and GPT-4V [33], thereby showing high transferability of adversarial examples across MLLMs. While extensive studies have explored adversarial attacks on image-based MLLMs, there has been little exploration in the domain of V-MLLMs. Li et al. [21] propose a flow-based adversarial attack strategy for white-box attacks on V-MLLMs. However, in real-world scenarios, the internal architectures and parameters of V-MLLMs are usually inaccessible to users. To address this, we focus on exploring methods for conducting adversarial attacks on V-MLLMs in a black-box setting.
2.3 Adversarial attack on video models
Current video models have diverse applications, including autonomous vehicles, video verification, security, and other fields. However, these models remain vulnerable to adversarial attacks [23, 40, 17, 38, 6]. For example, Universal 3D perturbations (U3D) [40] deceive video classifiers by generating a universal perturbation for all input videos, while StyleFool [5] introduces an unrestricted perturbation to attack video classification systems through style transfer. Recent studies also explore cross-modal attack methods from image models to video models [39, 19, 37], yielding promising results. However, these attacks primarily target video classification tasks, which do not account for interactions between visual and textual modalities. In contrast, V-MLLMs integrate both visual and textual information, rendering these methods unsuitable for such models. To address this limitation, our method incorporates multimodal interactions when crafting adversarial video samples, aligning with the operational principles of V-MLLMs.
@@ 3.1 Preliminary
Given a video sample $V\in\mathcal{V}$ with $M$ associated QA pairs $\{(q_{m},a_{m})\}_{m=1}^{M}$, where $q_{m}$ is the $m$-th question and $a_{m}$ is the corresponding answer. Let $F$ denote the IMM (e.g., BLIP-2 [20], MiniGPT-4 [46]) and $G$ denote the V-MLLM (e.g., Video-LLaMA [42], Chat-UniVi [18]). We use $G(V,q)$ to denote the answer generated by the V-MLLM for the given video $V$ and question $q$. The goal of our proposed attack is to generate an adversarial example $V_{adv}=V+\delta^{\prime}$ using $F$, which can cause $G$ to produce an answer $G(V_{adv},q_{i})$ that differs significantly from the correct answer $a_{i}$, without accessing the parameters or structure of $G$, where $\delta^{\prime}$ denotes the adversarial perturbations specifically tailored for $V$. To ensure that the adversarial perturbation $\delta^{\prime}$ is imperceptible, we restrict it by $||\delta^{\prime}||_{\infty}\leq\epsilon$, where $||\cdot||_{\infty}$ denotes the $L_{\infty}$ norm, and $\epsilon$ is a constant for the norm constraint. We utilize the evaluation model $E$ (i.e., GPT-4o-mini [32]) to assess whether the generated answer aligns with the reference answer. We aim to find adversarial perturbations that minimize the number of correct responses, formulated as follows:
$\underset{\delta^{\prime}}{\text{argmin}}\frac{1}{M}\sum_{i=1}^{M}E(G(V+\delta^{\prime},q_{i}),a_{i}),\text{s.t.}\ \ ||\delta^{\prime}||_{\infty}\leq\epsilon,$ (1)
@@
Attack Target Model
Chat-UniVi LLaVA-NeXT-Video VideoChat Video-LLaMA
FMM 8.11 15.38 14.62* 20.74
Vanilla 8.22 15.86 35.03* 19.81
I2V 25.17 27.39 27.13 30.51
FMM w/ Prop. 14.54 28.31 14.62* 27.38
Vanilla w/ Prop. 13.59 23.06 35.03* 27.98
I2V-MLLM 43.39 40.54 63.09 74.91
Table 1: Attack success rates (ASR, %) on the MSVD-QA validation set for VideoQA tasks. FMM and I2V denote attack methods from [21] and [39], respectively. Vanilla attack applies full perturbations on all key-frames sampled by V-MLLMs. Prop. denotes perturbation propagation. * indicates white-box attacks. A higher ASR indicates better adversarial transferability.
@@ 3.2 Motivation
Focusing only on sparse key-frames. The FMM attack exhibits limited transferability in the black-box setting due to differences in key-frame selection between the attack method and V-MLLMs. To address this, we first modify the FMM attack by replacing the sparse spatial perturbation with full perturbation on the key-frames sampled by V-MLLMs, which we call the Vanilla attack. While this adjustment improves white-box performance, the transferability still remains constrained. To further enhance transferability, we propagate perturbations from key-frames across the entire video, leading to improved transferability, as shown in rows 1, 2, 4, and 5 of Tab. 1.
Lacking generalization in perturbing video features. Adversarial perturbations generated based on certain V-MLLM can overfit to specific video module, limiting their generalization to other V-MLLMs. To improve transferability, we focus on lower-level image features. The I2V attack [39], which perturbs each video frame to disrupt image features, demonstrates improved transferability when using image models as surrogates to craft adversarial video samples, as shown in rows 3, 4, and 5 of Tab. 1.
Failing to integrate multimodal information. The I2V attack shows a limited improvement in transferability, as it was originally designed for video classification and does not account for the multimodal interactions, which is essential for V-MLLMs. Therefore, we propose using an image-based multimodal model as a surrogate, integrating multimodal interaction information into the process of generating adversarial video samples, which leads to a significant improvement in transferability, as demonstrated in rows 3 and 6 of Tab. 1.
In summary, we propose using IMMs as surrogates to generate adversarial video samples that incorporate multimodal interactions. In addition, we introduce a perturbation propagation technique to handle different unknown frame sampling strategies. The I2V-MLLM results in Tab. 1 demonstrate the strong transferability of our method across different V-MLLMs. More discussions can be found in Appendix A. The following sections describe our proposed attack in detail.

### v3

# Transferability of Adversarial Attacks in Video-based MLLMs: A Cross-modal Image-to-Video Approach

@@ Abstract
Video-based multimodal large language models (V-MLLMs) have shown vulnerability to adversarial examples in video-text multimodal tasks. However, the transferability of adversarial videos to unseen models—a common and practical real-world scenario—remains unexplored. In this paper, we pioneer an investigation into the transferability of adversarial video samples across V-MLLMs. We find that existing adversarial attack methods face significant limitations when applied in black-box settings for V-MLLMs, which we attribute to the following shortcomings: (1) lacking generalization in perturbing video features, (2) focusing only on sparse key-frames, and (3) failing to integrate multimodal information. To address these limitations and deepen the understanding of V-MLLM vulnerabilities in black-box scenarios, we introduce the Image-to-Video MLLM (I2V-MLLM) attack. In I2V-MLLM, we utilize an image-based multimodal large language model (I-MLLM) as a surrogate model to craft adversarial video samples. Multimodal interactions and spatiotemporal information are integrated to disrupt video representations within the latent space, improving adversarial transferability. Additionally, a perturbation propagation technique is introduced to handle different unknown frame sampling strategies. Experimental results demonstrate that our method can generate adversarial examples that exhibit strong transferability across different V-MLLMs on multiple video-text multimodal tasks. Compared to white-box attacks on these models, our black-box attacks (using BLIP-2 as a surrogate model) achieve competitive performance, with average attack success rate (AASR) of 57.98% on MSVD-QA and 58.26% on MSRVTT-QA for Zero-Shot VideoQA tasks, respectively.
Figure 1: An example of transferable adversarial attack on different target V-MLLMs for Zero-Shot VideoQA task.
Figure 2: Overview of our proposed method. (a) I2V-MLLM Attack: The clean video is divided into $K$ clips. Key frames are extracted from these clips to form the clean frames $X$, which is then fed into the vision model to extract clean frame-level embeddings $F_{V}(X)$. These embeddings are subsequently aggregated via spatiotemporal pooling to obtain clean spatiotemporal embeddings $F_{V}^{st}(X)$. Perturbations are initialized and added to clean frames $X$ to generate adversarial frames $X_{adv}$. The same process is applied to extract $F_{V}(X_{adv})$ and $F_{V}^{st}(X_{adv})$. An LLM reformulates the QA pairs into a caption set $T$. $F_{V}(X)$, $F_{V}(X_{adv})$, and $T$ are then passed through the projector to extract visual features $F_{P}^{v}(X)$, adversarial visual features $F_{P}^{v}(X_{adv})$, and textual features $F_{P}^{t}(T)$, respectively. Perturbations are updated via the PGD method by minimizing three cosine similarity-based losses: $L_{V}$, $L_{P}^{v}$, and $L_{P}^{v2t}$. (b) Perturbation Propagation: The final perturbations applied to key-frames are propagated back to their corresponding video clips to construct the adversarial video. (c) Attack Different Target V-MLLMs.
@@ 1 Introduction
Recent work has shown that video-based multimodal large language models (V-MLLMs) are vulnerable to adversarial video samples (li2024fmmattackflowbasedmultimodaladversarial), even though they have achieved remarkable performance on a wide range of video-text multimodal tasks (li2024videochatchatcentricvideounderstanding; jin2024chatuniviunifiedvisualrepresentation; lin2024videollavalearningunitedvisual; maaz2024videochatgptdetailedvideounderstanding; dai2023instructblipgeneralpurposevisionlanguagemodels; zhang2023videollamainstructiontunedaudiovisuallanguage). Existing work primarily focuses on white-box attacks, where information about the target model is accessible. However, the transferability of adversarial video samples across V-MLLMs remains unexplored, which is a more common and practical setting in real-world scenarios. It is still uncertain whether the adversarial videos generated on the source model can effectively attack other target models, posing significant security risks to the deployment of V-MLLMs in real-world applications.
@@
In this paper, we pioneer an investigation into the transferability of adversarial video samples across V-MLLMs. Through detailed analysis in Sec. 3.2, we think previous methods have these shortcomings: (1) lacking generalization in perturbing video features, (2) focusing only on sparse key-frames, and (3) failing to integrate multimodal information. FMM attack (li2024fmmattackflowbasedmultimodaladversarial) is the first proposed white-box attack method targeting V-MLLMs. It utilizes flow-based temporal mask to select key-frames and applies perturbations to these frames. FMM attack performs well in the white-box setting but has limited transferability in the black-box setting. FMM attack heavily relies on the video features, which causes the generated perturbations to overfit to the video features extracted by the surrogate model, thereby reducing their generalizability. Additionally, since FMM attack applies perturbations only to key-frames, it cannot ensure that all frames sampled by the target model are perturbed. Taking low-level image features into account can help with improving transferability of adversarial samples. Previous image-to-video cross-modal attacks (wei2021crossmodaltransferableadversarialattacks; wang2023global; kim2023breakingtemporalconsistencygenerating) demonstrate the possibility of using image models as surrogates to attack video models in the black-box setting. However, these traditional attack methods typically focus on the video classification tasks with vision-only models, failing to integrate multimodal information.
@@
To address these limitations, we propose a highly transferable attack method, named as Image To Video MLLM (I2V-MLLM) attack (see Fig. 2). In I2V-MLLM, we utilize an image-based multimodal large language model (I-MLLM) as a surrogate model to craft adversarial video samples without accessing the internals of target V-MLLMs. Specifically, we extract key-frames from videos and send them into an I-MLLM to obtain adversarial perturbations. Multimodal interactions and spatiotemporal information are integrated to disrupt video representations within the latent space, improving adversarial transferability. Additionally, perturbation propagation technique is introduced to handle different unknown frame sampling strategies used by V-MLLMs.
@@
We conduct various experiments on three well-established datasets, MSVD-QA (xu2017video), MSRVTT-QA (xu2017video), and ActivityNet-200 (Heilbron_Escorcia_Ghanem_Niebles_2015) to evaluate the performance of our proposed I2V-MLLM attack in multiple video-text multimodal tasks. The experimental results demonstrate that our method can generate adversarial videos with strong transferability across different V-MLLMs (Chat-Univi (jin2024chatuniviunifiedvisualrepresentation), LLava-Next-Video (zhang2024videoinstructiontuningsynthetic), VideoChat, Video-LLaMA (zhang2023videollamainstructiontunedaudiovisuallanguage)), and achieve competitive performance with white-box attacks against V-MLLMs. Our main contributions are summarized as follows:
@@
  • We explore the transferable adversarial attack on four V-MLLMs and analyze the reasons for the low transferability when using existing methods to generate adversarial video samples (see Sec. 3.2). To the best of our knowledge, this is the first work to explore black-box attacks on V-MLLMs.
@@
  • We propose a highly transferable attack method, named I2V-MLLM, for V-MLLMs using I-MLLMs to generate adversarial video samples (see Sec. 3.3). The adversarial videos generated by this method can effectively disrupt different V-MLLMs, significantly degrading their performance on multiple video-text multimodal tasks.
@@
  • We conduct extensive experiments on four different V-MLLMs using MSVD-QA, MSRVTT-QA, and ActivityNet-200 (see Sec. 4.2 and Appendix B). The results demonstrate that our proposed attack method has strong transferability across V-MLLMs.
@@ 2.1 Multimodal large language models
MLLMs typically consist of a vision model, a pretrained LLM, and a projector that translates visual information into textual representations that the LLM can process. Currently, MLLMs can be categorized into image-based and video-based types. I-MLLMs (dai2023instructblipgeneralpurposevisionlanguagemodels; liu2023visualinstructiontuning; zhu2023minigpt4enhancingvisionlanguageunderstanding; alayrac2022flamingovisuallanguagemodel; awadalla2023openflamingoopensourceframeworktraining; hu2023blivasimplemultimodalllm; bai2023qwenvlversatilevisionlanguagemodel) are designed to handle image-text inputs. They excel in tasks such as visual question answering, image captioning, and more. V-MLLMs extend the capabilities of I-MLLMs by incorporating temporal modules that allow them to understand and process video input. This enables them to perform tasks like video question answering (VideoQA), spatiotemporal localization, and video captioning. For example, Chat-UniVi (jin2024chatuniviunifiedvisualrepresentation) extracts specific frames from videos and utilizes DPC-KNN (Du_Ding_Jia_2016) to group these frames into distinct events, Video-LLaMA (zhang2023videollamainstructiontunedaudiovisuallanguage) employs sequential encoding to capture temporal relationships among video frames, VideoChatGPT (maaz2024videochatgptdetailedvideounderstanding) applies temporal pooling to video features to extract temporal information. These methods equip the models with the capability to capture and interpret temporal dynamics, thus enabling a more comprehensive understanding of video content.
@@ 2.2 Adversarial attacks on MLLMs
Despite the impressive performance, MLLMs are highly susceptible to adversarial attacks (zhao2023evaluatingadversarialrobustnesslarge; luo2024imageworth1000lies; cui2023robustnesslargemultimodalmodels; tu2023many; zhang2024avibench; bailey2023image; lu2023set). For I-MLLMs, several studies have assessed their vulnerabilities to adversarial attacks. (fu2023misusingtoolslargelanguage) introduces Trojan-like images that force the target models to invoke malicious external tools or APIs specified by the attacker. (dong2023robustgooglesbardadversarial) utilizes open-source MLLMs to generate transferable adversarial examples capable of attacking closed-source commercial models like Bard (google2023gemini), Bing Chat (microsoft2024bingchat), and GPT-4V (openai2024gpt4v), thereby showing high transferability of adversarial examples across MLLMs. (gu2025improving) introduces DynVLA Attack, which applies dynamic perturbations to the vision-language connector to improve the generalization of adversarial attacks across different alignment strategies. While extensive studies have explored adversarial attacks on I-MLLMs, there has been little exploration in the domain of V-MLLMs. (li2024fmmattackflowbasedmultimodaladversarial) proposes a flow-based adversarial attack strategy for white-box attacks on V-MLLMs. However, in real-world scenarios, the internal architectures and parameters of V-MLLMs are usually inaccessible to users. To address this, we focus on exploring methods for conducting adversarial attacks on V-MLLMs in a black-box setting.
2.3 Adversarial attacks on video models
Current video models have diverse applications, including autonomous vehicles, video verification, security, and other fields. However, these models remain vulnerable to adversarial attacks (li2018adversarial; xie2022universal; jiang2019black; wei2020heuristic; cao2024logostylefoolvitiatingvideorecognition). For example, Universal 3D perturbations (U3D) (xie2022universal) deceive video classifiers by generating a universal perturbation for all input videos, while StyleFool (cao2023stylefool) introduces an unrestricted perturbation to attack video classification systems through style transfer. Recent studies also explore cross-modal attack methods from image models to video models (wei2021crossmodaltransferableadversarialattacks; kim2023breakingtemporalconsistencygenerating; wang2023global), yielding promising results. However, these attacks primarily target video classification tasks, which do not account for interactions between visual and textual modalities. In contrast, V-MLLMs integrate both visual and textual information, rendering these methods unsuitable for such models. To tackle this issue, our method introduces multimodal interactions in adversarial video crafting, aligning seamlessly with how V-MLLMs operate.
@@ 3.1 Preliminary
Given a video sample $V\in\mathcal{V}$ with $M$ associated QA pairs $\{(q_{m},a_{m})\}_{m=1}^{M}$, where $q_{m}$ is the $m$-th question and $a_{m}$ is the corresponding answer. Let $F$ denote the I-MLLM (e.g., BLIP-2 (li2023blip2bootstrappinglanguageimagepretraining), MiniGPT-4 (zhu2023minigpt4enhancingvisionlanguageunderstanding)) and $G$ denote the V-MLLM (e.g., Video-LLaMA (zhang2023videollamainstructiontunedaudiovisuallanguage), Chat-UniVi. We use $G(V,q)$ to denote the answer generated by the V-MLLM for the given video $V$ and question $q$. The goal of our proposed attack is to generate an adversarial example $V_{adv}=V+\delta^{\prime}$ using $F$, which can cause $G$ to produce an answer $G(V_{adv},q_{i})$ that differs significantly from the correct answer $a_{i}$, without accessing the parameters or structure of $G$, where $\delta^{\prime}$ denotes the adversarial perturbations specifically tailored for $V$. To ensure that the adversarial perturbation $\delta^{\prime}$ is imperceptible, we restrict it by $\delta^{\prime}||_{\infty}\leq\epsilon$, where $||\cdot||_{\infty}$ denotes the $L_{\infty}$ norm, and $\epsilon$ is a constant for the norm constraint. We utilize the evaluation model $E$ (i.e., GPT-4o-mini (openai2023gpt4omini)) to assess whether the generated answer aligns with the reference answer. We aim to find imperceptible adversarial perturbations that minimize the number of correct responses, formulated as follows:
[…]
@@ 3.2 Motivation
Focusing only on sparse key-frames. FMM attack exhibits limited performance when the key-frame ratio is low (see Appendix A). This is because FMM selects key-frames based on optical flow and only perturbs these frames, while V-MLLMs typically sample frames sequentially, making it difficult to ensure that all frames extracted by the target model are perturbed. To address this, we first modify the FMM attack by replacing the sparse spatial perturbation with full perturbation on the key-frames sampled by V-MLLMs, which we call the Vanilla attack. While this adjustment improves white-box performance, the transferability still remains constrained. To further enhance transferability, we propagate perturbations from key-frames across the entire video, leading to improved transferability, as shown in rows 1, 2, 4, and 5 of Tab. 1.
Lacking generalization in perturbing video features. Adversarial perturbations generated based on certain V-MLLM can overfit to specific video module, limiting their generalization to other V-MLLMs. To improve transferability, we focus on lower-level image features. The I2V attack (wei2021crossmodaltransferableadversarialattacks), which perturbs each video frame to disrupt image features, demonstrates improved transferability when using image models as surrogates to craft adversarial video samples, as shown in rows 3, 4, and 5 of Tab. 1.
Failing to integrate multimodal information. The I2V attack shows a limited improvement in transferability, as it was originally designed for video classification and does not account for the multimodal interactions, which is essential for V-MLLMs. Therefore, we propose using an I-MLLM as a surrogate, integrating multimodal interaction information into the process of generating adversarial video samples, which leads to a significant improvement in transferability, as demonstrated in rows 3 and 6 of Tab. 1.
In summary, we propose using I-MLLMs as surrogates to generate adversarial videos that incorporate multimodal interactions. In addition, we introduce a perturbation propagation method to handle different unknown frame sampling strategies. The I2V-MLLM results in Tab. 1 demonstrate the strong transferability of our method across different V-MLLMs. More discussions can be found in Appendix A. The following sections detail the I2V-MLLM attack.
Attack Target Model AASR
Chat-UniVi LLaVA-NeXT-Video VideoChat Video-LLaMA
FMM 8.70 18.76 13.84 27.93* 17.31
Vanilla 11.62 25.31 15.10 64.14* 29.04
I2V 32.17 33.39 41.13 36.51 35.80
FMM w/ Prop. 14.54 25.31 14.62 27.93* 20.60
Vanilla w/ Prop. 14.94 32.24 17.50 64.14* 32.21
I2V-MLLM 48.39 45.54 63.09 74.91 57.98
Table 1: Attack success rates (ASR, %) on the MSVD-QA validation set for Zero-Shot VideoQA tasks. FMM and I2V denote attack methods from (li2024fmmattackflowbasedmultimodaladversarial) and (wei2021crossmodaltransferableadversarialattacks), respectively. Vanilla attack applies full perturbations on all key-frames sampled by V-MLLMs. Prop. denotes perturbation propagation. * indicates white-box attacks. AASR represents the average ASR across all target models for each surrogate model. A higher AASR indicates better adversarial transferability.